OpenStudy (lanhikari22):

Why do you have to write ln|x| when the domain simply can't refer to negative values?

OpenStudy (lanhikari22):

Also, since we do focus on that, why is it when doing calculations sometimes \[\sqrt(f(x)^2\] is taken to mean f(x) instead of |f(x)} ?

Parth (parthkohli):

so you mean why we don't write ln(x) instead of ln|x| well because in the second, the domain is different. x = -1 isn't in the domain of the first.

OpenStudy (lanhikari22):

In the second, sign information was removed by the squaring, so why did we not directly indicate that?

Parth (parthkohli):

also about your second question, it depends on the function. if x > 0 then sqrt(x^2) = x.

OpenStudy (misty1212):

HI!!

OpenStudy (dls):

\(\Large \ln(|x|)\) is defined on \(( - \infty, \infty)\) while while \(\Large \ln(x)\) is only defined on the interval \(( 0, \infty)\) If your x is defined on the former interval then your corresponding function you should use is ln (|x|) and vice versa.

OpenStudy (misty1212):

the domain of \(\ln(x)\) is \(x>0\) but the domain of \(\ln(|x|)\) is \(x\neq 0\)

OpenStudy (misty1212):

what @DLS said

OpenStudy (misty1212):

almost

OpenStudy (misty1212):

you still can't take the log of zero, so zero is out of both

OpenStudy (lanhikari22):

Oh, I see! So we're actually creating a new funtion that's a mirror of the right side!

OpenStudy (misty1212):

yes

OpenStudy (misty1212):

are you involved in calculus now?

OpenStudy (misty1212):

for your second question, sometimes it is assumed or arranged that \(f\) is always non-negative, so that \[\sqrt{f^2}=f\]

OpenStudy (lanhikari22):

Yeah.

OpenStudy (misty1212):

then you will know that the derivative of \(\ln(x)\) is \(\frac{1}{x}\) and also the derivative of \(\ln(-x)=\frac{1}{x}\) by the chain rule, so frequently people will write \[\int \frac{dx}{x}=\ln(|x|)\]

OpenStudy (misty1212):

as an example of \[\sqrt{f^2}=f\] sometimes they write \[\sqrt{\sin^2(x)}=\sin(x)\] but in those cases the domain is usually restricted to \[[0,\frac{\pi}{2}]\] where sine is positive

OpenStudy (dls):

Just like @misty1212 said, but do pay attention on the function you are dealing with though. For instance, \[\Large \sqrt{1 - 2\sin^2x} = |\cos x| \neq \cos x \]

OpenStudy (lanhikari22):

yes, it looks weird, not like a sin wave

OpenStudy (lanhikari22):

the downside is mirrored up

OpenStudy (lanhikari22):

but if \[\sqrt(\sin^2(x))\] would never be similar to sin(x), since |sin(x)| is a different function for ever half perod, how is that implicitly dealt with? What's the derivative of |sin x| ? when x is positive, it's simply cos(x), but does the function |sin x| still change according to that?